TRACE FORMS OF GALOIS FIELD EXTENSIONS IN THE PRESENCE OF ROOTS OF UNITY

D.-S. KANG ZINOVY REICHSTEIN

Abstract. Let L/K be a Galois field extension and let G_{2} be a Sylow subgroup of Gal(L/K). We show that if G_{2} is not abelian then the trace form of L/K is hyperbolic, provided that K contains certain roots of unity.